OpenStudy (anonymous):

factor: 100x^2-60x+9

OpenStudy (anonymous):

cheat

jhonyy9 (jhonyy9):

x1,2=(60+/- sqrt(3600-3600)/200 =(60+/- 0)/200 = 60/200 = 6/20 = 3/10 so this equation will be (10x -3)(10x-3)

jhonyy9 (jhonyy9):

or (10x-3) squared

OpenStudy (anonymous):

sir please help me it's my home work (x2+2x+3)(x2+4x+3)>0

jhonyy9 (jhonyy9):

(x2+2x+3)>0 and (x2+4x+3)>0 or (x2+2x+3)<0 and (x2+4x+3)<0 so than

jhonyy9 (jhonyy9):

x2+2x+3 is always gr

jhonyy9 (jhonyy9):

at

jhonyy9 (jhonyy9):

allways greater than zero

jhonyy9 (jhonyy9):

x2+4x+3 > 0 if x < -3 or x>-1 and have not possibility again thats all